OpenStudy (anonymous):

given that the slope of a line is -4/3 what is the slope of a line that is perpendicular to it

OpenStudy (sweetburger):

take the negative reciprocal of the slope to find the perpendicular slope

OpenStudy (anonymous):

would it then be 3/-4

OpenStudy (anonymous):

@sweetburger

OpenStudy (sweetburger):

you got the reciprocal part but you forgot to apply the negative

rishavraj (rishavraj):

if two line are mutually perpendicular and theire slopes are m1 and m2...then m1*m2 = -1

OpenStudy (anonymous):

im still confused so i got the reciprpcal but not the negative

OpenStudy (sweetburger):

alright you want to find the perpendicular slope to (-4/3) you took the reciprocal which is (-3/4) but forgot to do this -1(-3/4)

rishavraj (rishavraj):

m1*m2 = -1..here m1 = -4/3... m2 = -1/(-4/3)

OpenStudy (anonymous):

-1(-3/4) isnt an option though

rishavraj (rishavraj):

-1(-3/4) = 3/4......